OpenStudy (anonymous):

Scholastic Aptitude Test (SAT) scores are normally distributed with a mean of 500 points and a standard deviation of 100 points. Suppose you take the SAT and several weeks later you receive a letter telling you that your results on the math portion of the test were in the 95th percentile. Recalling that SAT scores are always expressed as multiples of 10, how many points did you get on the test?

OpenStudy (anonymous):

I really have no idea what to do or where to start here

OpenStudy (anonymous):

.95 * 500 = 475 / 10 = 47.5? not sure

OpenStudy (anonymous):

thats probs wrong

OpenStudy (anonymous):

hmm oh well ill look it over thanks

OpenStudy (amistre64):

(SAT) scores are normally distributed mean: 500 points sd: 100 points. Suppose you take the SAT and several weeks later you receive a letter telling you that your results on the math portion of the test were in the: 95th percentile. P{95} Recalling that SAT scores are always expressed as multiples of 10, how many points did you get on the test? Whats your z-score?

OpenStudy (amistre64):

z = (x - mean)/sd if i recall correctly; or just ti83 it :)

OpenStudy (amistre64):

P{95} = 664.4853...

OpenStudy (amistre64):

either 660 or 670

OpenStudy (amistre64):

x = z(sd) + mean .... but i aint got my z-table handy :) but it should get you to 1.64 1.64(100) = 164+500 = 664

OpenStudy (amistre64):

id say that if your "within" the P{95} that the score would be 660, but thats just a guess